convert 60.4 (the 4 has a line over it) to a fraction
Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(57214)

convert 60.4 (the 4 has a line over it) to a fraction
----
Let x = 60.444444...
Then 10x = 604.444444.....
-----
Subtract to get:
9x = 604-60 = 544
x = 544/9
----------------
Cheers,
Stan H.
=================

Answer by ewatrrr(10682)

 
Hi,
x = 60.4
10x = 10 ·60.4 = 604.4
10x -x = 604.4 -60.4 = 544
9x = 544
x = 544/9